# Quantifying the Impact of Incompleteness on Identifying and Interpreting Galaxy Protocluster Populations with the TNG-Cluster Simulation

### v1

@@ 3.3 Evaluating Completeness Limits for Recovering the True Highest Galaxy Density Peak
To compare galaxies in the highest matter density peaks to those outside these regions, we compute the ratio of average stellar mass and SFR inside and outside $R_{10}$, centered on these peaks. Fig. 13 shows the ratio of average SFR (top) and stellar mass (bottom) versus redshift for galaxies within $R_{10}$, centered on the highest dark matter density peak ($\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$), stellar mass density peak ($\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$), and galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$). Although not shown, we also examined the total baryonic mass density (stars + gas) and found it closely follows the dark matter density field. From left to right, the columns bin the ratios by protoclusters that will evolve into high-, intermediate-, and low-mass clusters by $z=0$. Fig. 13 presents the data as a box and whisker plot, with blue, orange, and yellow boxes showing the ratio distributions relative to $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$$\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$, and $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, respectively. Horizontal bars and circles indicate the median and mean of each distribution.
The top row of Fig. 13 shows that, whether probed by the mean or the median, the average SFR centered on $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$ and $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$ is significantly above unity for $z>1$. This contrasts sharply with the results relative to $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, where the median of the distribution is below unity for $z>2$. At $z=1$, all three distributions return comparable ratios of the average SFR within and beyond the respective highest density peak, reflecting the results from the top-left panel of Fig. 12, where the locations of these density peaks overlap by this epoch.
@@
Similarly, the bottom row of Fig. 13 shows that, whether assessed by the mean or the median, the typical stellar mass within and beyond $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$ and $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$ is also significantly above unity at $z\geq 2$, with earlier epochs showing the strongest enhancement. Once again, this contrasts sharply with the median of the distribution of the ratio of the average stellar mass within and beyond $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, which is below unity for $z>2$. However, by $z=1$ all three distributions return comparable ratios of the average stellar mass within and beyond the respective highest density peak.
The results in Fig. 13 show that without weighting by stellar mass, the highest density peaks traced by galaxy populations complete down to ${M}_{\star}=10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ do not generally correspond to regions of accelerated mass growth or enhanced star formation relative to the rest of the protocluster. This suggests that observed protocluster density maps, which are more heavily weighted towards more massive galaxies due to the incompleteness of lower-mass galaxies (e.g., ${M}_{\star}<10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) at $z>2$, may overestimate how often the inferred highest density regions are sites of accelerated galaxy evolution.
@@ 4 Discussion
4.1 Implications of Misidentifying the True Highest Galaxy Density Peak in Protoclusters
@@
The misidentification of the highest galaxy density peak has important consequences. When comparing galaxies near the highest galaxy density peak to those in the broader protocluster, we find that galaxies in these peaks have, on average, lower SFRs and stellar masses relative to those outside, seemingly contradicting previous studies that suggest the densest regions of protoclusters are sites of accelerated mass growth and enhanced star formation. This discrepancy arises because the highest galaxy density peaks do not necessarily coincide with the highest matter density peaks. Instead, the latter — traced by stellar mass and dark matter — exhibit the expected signs of rapid mass growth and high SFRs. Similarly, Muldrew et al. (2018) found that the largest progenitor halos in protoclusters host the highest stellar mass and SFR densities, attributing this to these halos being biased tracers of the underlying dark matter density field (Tinker et al., 2010). This suggests that observational studies, often limited to massive, star-forming protocluster galaxies, are prone to overestimating how frequently the inferred highest-density regions are sites of accelerated stellar mass growth and enhanced star formation, as they likely miss the true highest-density regions.
If future observations probe lower-mass galaxies (e.g., ${M}_{\star}\gtrsim 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$), they may find that the highest galaxy density regions are not necessarily sites of enhanced galaxy evolution compared to the broader protocluster population. However, it is unclear whether these galaxies will exhibit enhanced SFRs and accelerated mass growth compared to a coeval field population. This would be an interesting avenue for exploration, as current observational evidence suggests that while the densest regions in protoclusters show star formation enhancement relative to the field, the majority of protocluster members are on the main sequence (Shimakawa et al., 2018b; Polletta et al., 2021; Pérez-Martínez et al., 2023, 2024a, 2024b). However, Hayashi et al. (2016) found that relatively low-mass galaxies (${M}_{\star}\gtrsim 10^{9.3}\leavevmode\nobreak\ {\rm M}_{\odot}$) in the USS 1558-003 protocluster at $z=2.5$ exhibit higher SFRs than main-sequence galaxies at this redshift. These results highlight the need to expand the stellar mass and SFR limits of observational samples to determine if these trends extend to unseen members and to acquire large, spectroscopically-confirmed protocluster samples to assess how variance in protocluster evolutionary states influences these trends.
4.2 Considerations for Observational Studies of Galaxy Protoclusters
Our analysis suggests a few key considerations for observational studies of galaxy protoclusters. First, aperture sizes smaller than $4\arcmin$ should be avoided, as they fail to capture the full extent of the progenitors of the most massive galaxy clusters. These apertures also probe a region of confusion, as they can overlap with the innermost regions of the progenitors of the most massive clusters and the full extent of the progenitors of the lowest mass clusters (see Fig. 2). Instead, apertures larger than $8\arcmin$ are recommended, as they will capture the full extent of the progenitors of the most massive clusters at $z>2$.
@@ 4.2 Considerations for Observational Studies of Galaxy Protoclusters
Finally, the incompleteness in current protocluster selection functions, both in terms of stellar mass and SFR, inherently biases observed density maps. As shown in Fig. 13, the identification of the highest galaxy density peaks as regions of enhanced star formation and stellar mass growth depends on how the maps are weighted. While it is difficult to fully correct for this missing data, one approach would be to analyze analogs of the observed protoclusters at similar redshifts in cosmological simulations. Information provided by these models, such as the abundance, spatial distribution, and clustering of galaxies below the SFR and stellar mass sensitivity limits of the observations, could inform adjustments to density maps.
@@ 5 Summary and Conclusions
In this paper we leverage the unique combination of a large sample of very massive galaxy clusters (and their progenitors) and high baryonic mass resolution of the TNG-Cluster simulation to examine how observational incompleteness impacts inferences about galaxy protocluster populations. We achieve this by first defining our protocluster galaxy population as the ensemble of galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ by $z=0$. We define the inner and outer extents of galaxy protoclusters using $R_{10}$ and $R_{90}$, which represent the radii, measured relative to the center of mass, that enclose $10\%$ and $90\%$ of the total stellar mass traced by the baseline protocluster population. Lastly, we create density maps using adaptive binning weighted by nearest neighbor distance, which is qualitatively similar to results from a Kernel Density Estimation approach. With this information we investigate how the typical angular sizes used to define spectroscopically-confirmed protocluster samples compare to theoretical expectations, the impact of stellar mass and SFR incompleteness on identifying the highest galaxy density peak in protoclusters, and the relationship between local galaxy density variations in protoclusters and average galaxy properties. Our main conclusions are as follows:
@@
    The field of view of observational surveys typically used to define spectroscopically-confirmed protoclusters (about $\sim 4\arcmin$) are generally much smaller than the $8\arcmin$ apertures needed to study the progenitors of the most massive galaxy clusters at $3\leq z\leq 5$.
@@
    Angular sizes of $4\arcmin$, which are typical in defining spectroscopically-confirmed protocluster, fall within a region of confusion where they could be probing either the innermost region of a massive cluster progenitor or the full extent of a low-mass cluster progenitor.
@@
    Stellar mass-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR-limited (SFR$>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr^{-1}}$) protocluster populations recover only $\lesssim 40\%$ of the true highest galaxy density peaks within an accuracy of 1 pMpc ($\sim 2-2.5\leavevmode\nobreak\ \arcmin$) at $3\leq z\leq 5$, indicating that the bulk of the densest regions probed by observationally-limited protocluster samples are misidentified.
  4. (iv)
    For protocluster populations complete down to ${M}_{\star}\sim 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, inferences about whether the highest galaxy density peak corresponds to a region of accelerated stellar mass growth and enhanced star formation are sensitive to how the density map is weighted. Density maps that are agnostic to stellar mass typically identify highest galaxy density peaks that do not generally correspond to regions of accelerated galaxy evolution, whereas applying stellar mass weighting shifts these peaks toward regions of enhanced star formation and stellar mass growth. This disconnect likely arises because, at $3\leq z\leq 5$, the highest galaxy density peaks do not generally coincide with the highest matter density peaks.
  5. (v)
    Our results suggest that observed protocluster density maps, which are inherently more heavily weighted toward more massive galaxies due to the incompleteness of lower-mass galaxies (e.g., ${M}_{\star}<10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) at $z>2$, may overestimate the frequency with which the highest density regions are sites of accelerated galaxy evolution relative to the rest of the protocluster.
@@
This investigation focused on the stellar mass and SFR limitations of protocluster populations, but future studies could examine the impact of incompleteness in other observable quantities, such as luminosity and color, which can be modeled with simulations like TNG-Cluster. Additionally, simulation studies that begin with the observer’s definition of protoclusters as high-redshift overdensities — without a priori knowledge of their eventual evolution into clusters — will help elucidate how selection functions and incompleteness influence the fate of high-redshift overdensities identified in observations. Furthermore, detailed comparisons using mock observations, accounting for uncertainties in protocluster membership and selection functions, will be helpful in refining our interpretation of observed protocluster samples and prepare for future surveys (e.g., Euclid Collaboration et al., 2025b). These efforts are timely, as combining the ultra-deep imaging, wide fields of view, and high-resolution spectroscopy from recently launched and next-generation observatories such as JWST, LSST, Euclid, and Roman will lead to an explosion in the number of spectroscopically-confirmed and heterogeneously-selected protocluster samples, significantly enhancing our window into the earliest stages of galaxy cluster formation.
@@
DCB is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303800. DCB is also supported by the UC Chancellor’s Postdoctoral Fellowship. ALC is supported by the Ingrid and Joseph W. Hibben endowed chair at UC San Diego. DN acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) through an Emmy Noether Research Group (grant number NE 2441/1-1). AP acknowledges funding from the European Union (ERC, COSMIC-KEY, 101087822, PI: Pillepich) The TNG-Cluster simulation has been executed on several machines: with compute time awarded under the TNG-Cluster project on the HoreKa supercomputer, funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research; the bwForCluster Helix supercomputer, supported by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG; the Vera cluster of the Max Planck Institute for Astronomy (MPIA), as well as the Cobra and Raven clusters, all three operated by the Max Planck Computational Data Facility (MPCDF); and the BinAC cluster, supported by the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 37/935-1 FUGG. Many thanks to Maria Polleta for providing feedback that has greatly enhanced the quality of this manuscript. This research made extensive use of Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013a, 2018a). Additionally, the Python packages NumPy (Van Der Walt et al., 2011), iPython (Pérez & Granger, 2007), SciPy (Virtanen et al., 2020), and matplotlib (Hunter, 2007) were utilized for our data analysis and presentation. In addition, this research has made use of NASA’s Astrophysics Data System Bibliographic Services.

### v2

@@ 3.3 Evaluating Completeness Limits for Recovering the True Highest Galaxy Density Peak
To compare galaxies in the highest matter density peaks to those outside these regions, we compute the ratio of average stellar mass and SFR inside and outside $R_{10}$, centered on these peaks. Once again, to ensure a consistent comparison, we continue to measure $R_{10}$ relative to the center of mass of the baseline protocluster protocluster, as described in §2.4. Fig. 13 shows the ratio of average SFR (top) and stellar mass (bottom) versus redshift for galaxies within $R_{10}$, centered on the highest dark matter density peak ($\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$), stellar mass density peak ($\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$), and galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$). Although not shown, we also examined the total baryonic mass density (stars + gas) and found it closely follows the dark matter density field. From left to right, the columns bin the ratios by protoclusters that will evolve into high-, intermediate-, and low-mass clusters by $z=0$. Fig. 13 presents the data as a box and whisker plot, with blue, orange, and yellow boxes showing the ratio distributions relative to $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$$\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$, and $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, respectively. Horizontal bars and circles indicate the median and mean of each distribution.
The top row of Fig. 13 shows that, whether probed by the mean or the median, the average SFR ratios measured relative to $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$ , $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$ , and $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$  are above unity for $z>1$, implying that these are all regions of enhanced star formation activity relative to the rest of the protocluster. However, these ratios are relatively higher for the $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$  and $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$.
@@
Similarly, the bottom row of Fig. 13 shows that, whether assessed by the mean or the median, the average stellar mass within and beyond $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$ , $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$ , and $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$  is also significantly above unity at all redshifts. Once again, the ratios centered on $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$  and $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$are relatively higher than the ratios centered on $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$.
@@ 4 Discussion
4.1 How often is the Protocluster Core Misidentified?
In Chiang et al. (2017), a theoretical picture of the three stages of cluster assembly is presented, beginning with the inside-out growth of a protocluster “core”, followed by a phase of widespread star formation across the entire structure, and concluding with collapse and quenching. Within this framework, there is a disconnect between how the core of protoclusters is defined observationally and in simulations. For instance, in observations the core is typically identified as the region of highest galaxy density within the aperture used to detect the protocluster. However, in simulations the core is more often defined as the largest progenitor halo within the protocluster (e.g., Muldrew et al., 2018) or the most massive halo at a given epoch (Chiang et al., 2017; Araya-Araya et al., 2024). Following the observational definition, in this study we consider the region of highest galaxy density in our simulated protocluster sample as a proxy for protocluster core.
@@ 4.1 How often is the Protocluster Core Misidentified?
4.2 The Impact of Misidentifying the Protocluster Core
Observational protocluster studies show that the region of highest galaxy density – which again is often interpreted as the protocluster core – is generally associated with enhanced star formation activity and stellar mass growth. Similar trends have been noted in simulations, for instance Muldrew et al. (2018) found that the largest progenitor halos in protoclusters host the highest stellar mass and SFR densities, attributing this to these halos being biased tracers of the underlying dark matter density field. In this study we have demonstrated that these definitions for the protocluster core are generally inconsistent, with the location of the highest galaxy density peak and BCG progenitor agreeing in $\lesssim 45-55\%$ of cases within an accuracy of $1.0$ pMpc (see Fig. 8).
Nevertheless, we find that conclusions regarding the densest regions as sites of enhanced stellar mass growth and star formation activity are largely agnostic to how the protocluster core is defined. We also show that this conclusion remains unchanged even when the location of the highest-density peaks is determined from protocluster populations with stellar mass completeness limits offset by $1.5$ dex. However, the magnitude of the enhancement is somewhat higher when the peak is identified using a more complete protocluster population. These findings suggest that, if other sources of uncertainty — such as projection effects and redshift-space distortions — can be accounted for, then inferences about environment-driven enhancement in protoclusters are likely to be robust against variations in stellar mass completeness.
4.3 Link Between Galaxy and Matter Density Peaks
@@
4.4 Impact of Aperture Size on Protocluster Identification and Characterization
As highlighted in Fig. 3, our results suggest that radial aperture sizes smaller than $4\arcmin$ should be avoided, as they fail to capture regions beyond the inner core of galaxy protoclusters, regardless of the $z=0$ halo mass. To capture the full extent of the progenitors of the most massive clusters at $z>2$, aperture sizes greater than $8\arcmin$ are recommended. Using consistent and appropriate aperture sizes is crucial because they significantly impact the estimation of the $z=0$ halo mass (e.g., Lim et al., 2024), which is one approach used in the literature to determine whether a high-redshift overdensity is a true cluster progenitor. Aperture sizes can also impact estimates of the relative contribution of protoclusters to the cosmic star formation rate density (e.g., Chiang et al., 2017). Lastly, without exploring the full expected volume, the protocluster core may be misidentified as a local rather than the global highest galaxy density peak.
4.5 BCG Progenitor Identification
In Fig. 8 we show that the most massive protocluster galaxies coincide with the highest galaxy density peaks in $\lesssim 40\%$ of cases, within an accuracy of $1.0$ pMpc at $z>2$. However, the BCG progenitors — which are more challenging to identify observationally — are located within $1.0$ pMpc of the highest galaxy density peak in $\lesssim 60\%$ of cases at $z>2$. Contemporary datasets have enabled efforts to identify BCG progenitors in protocluster environments, as demonstrated in recent work by Shi et al. (2024) and Ito et al. (2025). The results from our analysis suggest that a potential effective observational strategy to accurately identify BCG progenitor candidates is to restrict the selection to galaxies that are both among the most massive in the structure and located within $1.0$ pMpc of the highest density peak. From a theoretical perspective, this is well motivated, as BCGs likely form through multiple mergers that occur preferentially in the densest region of the protocluster.
@@ 5 Summary and Conclusions
In this paper we leverage the unique combination of a large sample of very massive galaxy clusters (and their progenitors) and high baryonic mass resolution of the TNG-Cluster simulation to examine how observational incompleteness impacts inferences about galaxy protocluster populations. We achieve this by first defining our baseline protocluster galaxy population as the ensemble of galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that are progenitors of the galaxies that reside within $R_{\rm{200c}}$ at $z=0$. We define the inner and outer extents of galaxy protoclusters using $R_{10}$ and $R_{90}$, which represent the radii, measured relative to the center of mass, that enclose $10\%$ and $90\%$ of the total stellar mass traced by the baseline protocluster population. Lastly, we create density maps using adaptive binning weighted by nearest neighbor distance, which is qualitatively similar to results from a Kernel Density Estimation approach. With this information we investigate how the typical angular sizes used to define spectroscopically confirmed protocluster samples compare to theoretical expectations, the impact of stellar mass and SFR incompleteness on identifying the highest galaxy density peak in protoclusters (adopted here as a proxy for the protocluster core), and the relationship between local galaxy density variations in protoclusters and average galaxy properties. Our main conclusions are as follows:
@@
    The definition of the protocluster core is strongly influenced by the choice of aperture size: Typical radial apertures used in observational surveys ($\sim 4\arcmin$) are significantly smaller than the $\sim 8\arcmin$ apertures needed to capture the full extent of the progenitors of the most massive clusters at $3\leq z\leq 5$. As a result, such limited apertures may capture only the inner regions of protoclusters, leading to core definitions based on local rather than global galaxy density peaks.
@@
    The definition of the protocluster core is sensitive to stellar mass completeness: While stellar mass-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR-limited (SFR$>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr^{-1}}$) subpopulations recover the true highest galaxy density peak in $\lesssim 65\%$ of cases within $1.0$ pMpc ($2.1-2.6\arcmin$) at $3\leq z\leq 5$, more commonly-used higher mass cuts (e.g., ${M}_{\star}>10^{10}\leavevmode\nobreak\ {\rm M}_{\odot}$) reduce this recovery fraction to $50\%$. This suggests that typical observational completeness limits can lead to the misidentification of protocluster cores in a substantial fraction of cases.
@@
    Inferences about accelerated galaxy evolution in the protocluster core vary based on how the core is defined: Consistent with observations, the highest galaxy density peaks identified in our baseline protocluster population exhibit enhanced star formation activity and stellar mass growth compared to the rest of the protocluster. Similar signs of enhancements are measured when the highest density peaks are identified using a subpopulation with ${M}_{\star}>10^{10}\leavevmode\nobreak\ {\rm M}_{\odot}$, though the level of the enhancements are systematically lower than those seen in the baseline population.
@@
This investigation focused on the stellar mass and SFR limitations of protocluster populations; future studies could examine the impact of incompleteness in other observable quantities such as luminosity and color, which can be modeled with simulations like TNG-Cluster. There is also a need to explore the ramifications of dust extinction in constructing ${M}_{\star}$- and SFR-limited populations. Additionally, simulation studies that begin with the observer’s definition of protoclusters as high-redshift overdensities — without a priori knowledge of their eventual evolution into clusters — will help elucidate how selection functions and incompleteness influence the fate of high-redshift overdensities identified in observations. Furthermore, detailed comparisons based on mock observations using lightcones — accounting for uncertainties in protocluster membership and selection functions — will be valuable for refining our interpretation of observed protocluster samples and preparing for future surveys (e.g., Euclid Collaboration et al., 2025b). These efforts are timely, as combining the ultra-deep imaging, wide fields of view, and high-resolution spectroscopy from recently launched and next-generation observatories such as JWST, LSST, Euclid, and Roman will lead to an explosion in the number of spectroscopically confirmed and heterogeneously-selected protocluster samples, substantially enhancing our window into the earliest stages of galaxy cluster formation.
Acknowledgments
DCB is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303800. DCB is also supported by the UC Chancellor’s Postdoctoral Fellowship. ALC is supported by the Ingrid and Joseph W. Hibben endowed chair at UC San Diego. DN acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) through an Emmy Noether Research Group (grant number NE 2441/1-1). AP acknowledges funding from the European Union (ERC, COSMIC-KEY, 101087822, PI: Pillepich).
@@ Acknowledgments
The TNG-Cluster simulation has been executed on several machines: with compute time awarded under the TNG-Cluster project on the HoreKa supercomputer, funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research; the bwForCluster Helix supercomputer, supported by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG; the Vera cluster of the Max Planck Institute for Astronomy (MPIA), as well as the Cobra and Raven clusters, all three operated by the Max Planck Computational Data Facility (MPCDF); and the BinAC cluster, supported by the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 37/935-1 FUGG.
We thank Maria Polleta for providing feedback that has greatly enhanced the quality of this manuscript. We also thank the anonymous reviewer for their thoughtful comments that have enhanced the quality and clarity of this manuscript.
This research made extensive use of Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013a, 2018a). Additionally, the Python packages NumPy (Van Der Walt et al., 2011), iPython (Pérez & Granger, 2007), SciPy (Virtanen et al., 2020), and matplotlib (Hunter, 2007) were utilized for our data analysis and presentation. In addition, this research has made use of NASA’s Astrophysics Data System Bibliographic Services.